OpenStudy (anonymous):

Find g(–2) if g(x) = 4x + 3.

OpenStudy (anonymous):

g(-2) = 4*-2 + 3 = -8 + 3 = -5

Directrix (directrix):

4(-2) + 3 = -8 + 3 = -5

OpenStudy (anonymous):

g(-2 ) = 4(-2) + 3 = -5